# ShadowKV: KV Cache in Shadows for High-Throughput Long-Context LLM Inference

### v1

@@ 4.2 Theoretical Equivalent Bandwidth
The benefit of ShadowKV in terms of increasing throughput can be analyzed through the concept of equivalent bandwidth. Consider each K or V vector as being bytes in size, with a sequence length of , a chunk size of , a selected chunk budget of , outliers, and hit rate . During KV selection, ShadowKV loads approximately bytes using the GPU memory bandwidth . For value cache fetching, it loads bytes using the PCIe bandwidth [42]. Since value movement and key cache reconstruction can be overlapped, we do not need to count key cache reconstruction here. Following this, ShadowKV performs standard attention computation for the top-k chunks and predefined outliers, requiring bytes. The equivalent bandwidth of ShadowKV can be defined as:
For example, assuming C=8, S=128K, K=256, O=48, =31.5 GB/s, and =2 TB/s for A100, the equivalent bandwidth of ShadowKV is calculated as 7.2 TB/s, which is 3.6 higher than A100 memory bandwidth. This result indicates that ShadowKV theoretically achieves a high equivalent bandwidth to accelerate attention computation. System implementation is detailed in Section A.1.
Methods | N-S1 | N-S2 | N-MK1 | N-MK2 | N-MQ | N-MV | QA-1 | QA-2 | VT | FWE | Avg. |
Llama-3-8B-1M | 100.00 | 100.00 | 98.96 | 98.96 | 98.96 | 95.57 | 75.00 | 48.96 | 78.54 | 71.85 | 86.68 |
Loki | 18.75 | 1.04 | 2.08 | 0.00 | 1.56 | 0.78 | 4.17 | 13.54 | 26.04 | 25.35 | 9.33 |
Loki (V only) | 41.67 | 6.25 | 37.50 | 1.04 | 8.07 | 30.73 | 10.42 | 19.79 | 51.67 | 37.50 | 24.46 |
Quest | 100.00 | 100.00 | 98.96 | 77.08 | 97.65 | 93.49 | 60.42 | 50.00 | 77.08 | 65.63 | 82.03 |
Quest (V only) | 100.00 | 100.00 | 98.96 | 85.42 | 97.92 | 95.49 | 70.83 | 46.88 | 78.75 | 65.63 | 83.99 |
ShadowKV | 100.00 | 100.00 | 97.92 | 98.96 | 96.88 | 95.83 | 72.92 | 52.08 | 81.67 | 72.57 | 86.88 |
GLM-4-9B-1M | 100.00 | 100.00 | 94.79 | 87.50 | 99.74 | 93.75 | 67.71 | 55.21 | 97.29 | 72.22 | 86.82 |
Loki | 71.88 | 27.08 | 22.92 | 2.08 | 9.90 | 11.46 | 28.13 | 27.08 | 31.04 | 54.17 | 28.57 |
Loki (V only) | 96.88 | 55.21 | 56.25 | 18.75 | 51.04 | 50.52 | 45.83 | 39.58 | 72.71 | 59.72 | 54.65 |
Quest | 100.00 | 95.83 | 90.62 | 54.17 | 94.01 | 76.30 | 55.21 | 52.08 | 95.83 | 64.58 | 77.86 |
Quest (V only) | 100.00 | 96.88 | 93.75 | 72.92 | 95.83 | 83.07 | 56.25 | 53.13 | 96.88 | 65.97 | 81.47 |
ShadowKV | 100.00 | 100.00 | 95.83 | 83.33 | 98.70 | 87.76 | 69.79 | 55.21 | 97.50 | 68.06 | 85.62 |
Llama-3.1-8B | 100.00 | 100.00 | 98.96 | 91.67 | 98.96 | 95.31 | 82.29 | 47.92 | 68.96 | 71.18 | 85.53 |
Loki | 68.75 | 32.29 | 32.29 | 20.83 | 42.71 | 28.65 | 41.67 | 33.33 | 24.79 | 29.86 | 35.52 |
Loki (V only) | 95.83 | 36.46 | 57.29 | 62.50 | 77.86 | 70.83 | 69.79 | 39.58 | 35.21 | 37.50 | 58.29 |
Quest | 100.00 | 98.96 | 97.92 | 34.38 | 93.49 | 88.54 | 70.83 | 44.79 | 65.63 | 68.40 | 76.29 |
Quest (V only) | 100.00 | 98.96 | 98.96 | 56.25 | 95.83 | 90.63 | 76.04 | 46.88 | 66.25 | 67.36 | 79.72 |
ShadowKV | 100.00 | 100.00 | 100.00 | 83.33 | 97.92 | 92.19 | 81.25 | 48.96 | 67.08 | 64.93 | 83.57 |
Yi-9B-200K | 100.00 | 100.00 | 86.46 | 62.50 | 64.58 | 32.55 | 44.79 | 39.58 | 36.87 | 89.93 | 65.73 |
Loki | 34.38 | 2.08 | 2.08 | 0.00 | 0.00 | 0.52 | 22.92 | 21.88 | 0.00 | 25.00 | 10.89 |
Loki (V only) | 59.38 | 11.46 | 18.75 | 5.21 | 4.43 | 2.08 | 22.92 | 31.25 | 0.00 | 35.07 | 19.06 |
Quest | 100.00 | 98.96 | 79.17 | 26.04 | 56.51 | 31.77 | 32.29 | 31.25 | 51.04 | 71.88 | 57.89 |
Quest (V only) | 100.00 | 100.00 | 80.21 | 45.83 | 59.37 | 31.90 | 36.45 | 34.37 | 53.54 | 71.88 | 61.36 |
ShadowKV | 100.00 | 100.00 | 82.29 | 67.71 | 63.28 | 31.51 | 43.75 | 38.54 | 56.04 | 72.22 | 65.53 |
@@ 5 Empirical Evaluation
In Section 5.2, we demonstrate ShadowKV can support up to larger batch sizes and increase the inference throughput by up to without compromising model quality.
In Section 5.3, we present extensive ablation studies that validate the effectiveness of each component of ShadowKV in optimizing GPU memory usage and enhancing performance.
All details (hyperparameters, datasets, etc.), along with additional experiments, are in Appendix A.
@@ 5.1 Accuracy Evaluation
We demonstrate that ShadowKV can reduce the GPU memory usage of the KV cache by while maintaining accuracy on a range of long-context tasks with a minimal sparse KV cache budget.
@@
We choose four widely used long-context models for our evaluation: Llama-3-8B-1M [15], GLM-4-9B-1M [14], Llama-3.1-8B [33], and Yi-9B-200K [3]. We evaluate our approach on three challenging long-context benchmarks: RULER [20], LongBench [4], and Needle In A Haystack [24], covering QA, multi-hop, reasoning, summarization, code completion. Needle In A Haystack is also tested on Phi-3-Mini-128K [1] and Qwen2-7B-128K [59]. We set the chunk size to 8, the rank to 160, and the number of outliers to 48 for ShadowKV.
@@
We include two dynamic sparse attention methods as baselines: Quest [48] and Loki [43]. For all methods, we retain exact pre-filling and perform dynamic sparse attention during decoding, where the computation cost is set to 1/16 of full attention for selecting sparse KV pairs. We include two variants for each baseline: one where all the KV cache is offloaded, and another where only the value cache is offloaded. The former has similar latency to ShadowKV but a smaller sparse budget since ShadowKV only needs to fetch the value cache from the CPU. The latter aligns with the same sparse KV cache budget but significantly increases GPU memory usage. The latter one is marked as “V only” in the table.
Methods | NarratQA | MultiFQA | HotpotQA | MuSiQue | DuRead | GovRep | SAMSum | PassRetr | LCC | Avg. |
Llama-3-8B-1M | 18.98 | 41.84 | 36.79 | 21.47 | 31.93 | 34.18 | 35.96 | 81.50 | 56.07 | 39.86 |
Loki | 2.26 | 10.19 | 5.48 | 3.16 | 12.17 | 28.97 | 7.84 | 40.52 | 31.44 | 15.78 |
Loki (V only) | 3.20 | 21.01 | 12.41 | 3.86 | 17.07 | 31.24 | 16.23 | 52.57 | 38.10 | 21.74 |
Quest | 20.13 | 36.63 | 35.00 | 18.14 | 24.55 | 27.11 | 35.63 | 79.00 | 53.64 | 36.65 |
Quest (V only) | 17.26 | 39.51 | 36.78 | 18.71 | 26.41 | 29.49 | 35.80 | 79.50 | 60.05 | 38.17 |
ShadowKV | 17.17 | 39.73 | 38.29 | 21.08 | 31.77 | 31.62 | 35.87 | 80.00 | 63.93 | 39.94 |
GLM-4-9B-1M | 25.44 | 51.09 | 58.67 | 39.61 | 32.04 | 29.97 | 40.31 | 99.00 | 58.02 | 48.24 |
Loki | 5.82 | 30.60 | 22.73 | 9.20 | 30.09 | 30.35 | 22.70 | 98.92 | 40.77 | 32.35 |
Loki (V only) | 10.89 | 44.97 | 45.44 | 23.51 | 32.07 | 30.56 | 35.34 | 99.50 | 50.27 | 41.39 |
Quest | 23.81 | 44.53 | 56.41 | 35.49 | 23.54 | 21.73 | 37.39 | 87.00 | 43.80 | 41.52 |
Quest (V only) | 26.00 | 46.32 | 57.54 | 36.42 | 24.58 | 24.52 | 37.71 | 93.50 | 46.52 | 43.68 |
ShadowKV | 26.50 | 51.31 | 59.09 | 38.87 | 32.92 | 28.54 | 38.70 | 96.50 | 58.55 | 47.89 |
Llama-3.1-8B | 31.56 | 55.10 | 57.65 | 29.46 | 35.26 | 34.45 | 29.84 | 100.00 | 67.31 | 48.96 |
Loki | 2.31 | 18.89 | 10.64 | 5.47 | 19.30 | 31.16 | 15.91 | 94.88 | 44.60 | 27.02 |
Loki (V only) | 3.93 | 38.59 | 22.85 | 12.96 | 27.43 | 32.22 | 26.43 | 98.25 | 56.11 | 35.42 |
Quest | 29.70 | 49.04 | 53.96 | 27.18 | 27.16 | 30.43 | 29.85 | 98.50 | 57.35 | 44.80 |
Quest (V only) | 30.02 | 53.97 | 56.39 | 27.06 | 29.06 | 31.65 | 30.23 | 99.00 | 63.89 | 46.81 |
ShadowKV | 30.93 | 55.20 | 57.32 | 29.13 | 31.85 | 32.79 | 30.40 | 99.50 | 66.03 | 48.13 |
Yi-9B-200K | 13.88 | 30.02 | 52.46 | 28.20 | 22.29 | 30.25 | 19.08 | 67.00 | 73.50 | 37.41 |
Loki | 1.63 | 2.73 | 16.21 | 4.87 | 4.75 | 2.13 | 4.95 | 0.00 | 38.72 | 8.44 |
Loki (V only) | 1.96 | 10.39 | 21.31 | 7.36 | 6.78 | 9.15 | 10.02 | 4.00 | 58.75 | 14.41 |
Quest | 10.57 | 25.83 | 46.06 | 23.04 | 17.09 | 17.11 | 20.59 | 50.50 | 67.70 | 30.94 |
Quest (V only) | 14.56 | 25.73 | 48.73 | 24.73 | 18.44 | 20.83 | 20.08 | 57.50 | 71.13 | 33.53 |
ShadowKV | 12.44 | 30.82 | 52.43 | 27.73 | 20.79 | 29.83 | 20.73 | 64.00 | 72.89 | 36.85 |
@@
As shown in Table 1, ShadowKV demonstrates excellent performance on 128K contexts. With a fixed sparse budget of 1.56%, Loki and Quest experience performance degradation. In contrast, ShadowKV is more robust and even outperforms original full attention on certain tasks, such as variable tracking. For complex tasks like multi-document QA or multi-key needle retrieval, other methods suffer from significant performance degradation while ShadowKV does not.
@@
On LongBench, we evaluate our method with a range of realistic scenarios, including single-/multi-document question-answering, document summarization, code completion, information retrieval, etc. We only test on samples longer than 4K and set the sparse KV cache budget to 256 for this benchmark since it has shorter inputs compared to RULER. As shown in Table 2, ShadowKV outperforms other methods consistently and maintains the performance.
Needle In A Haystack.
On the Needle In A Haystack dataset, as shown in Figure 6, ShadowKV shows the ability to process information at different positions across various context windows, ranging from 16K to 1M tokens. More experiments on a range of models can be found in Section A.3.
@@
Llama-3-8B-1M w/ MInference | 89.92 | 88.02 | 82.81 | 78.45 | 78.12 | 74.57 | 81.98 |
ShadowKV w/ MInference | 90.47 | 88.12 | 83.28 | 77.71 | 78.32 | 74.31 | 82.04 |
Multi-turn Conversation Capability.
As the input context length scales, pre-filling becomes quite costly due to the quadratic growing computation time for attention, which means multi-turn conversation capability is important. To simulate multi-turn conversations, we challenged ShadowKV with a multi-turn needle retrieval task (Multi-turn NIAH). We also test two eviction-based methods in Figure 7, including SnapKV [29] and StreamingLLM [57]. The performance of SnapKV drops significantly from the second round due to the required context information being different from the first round. Since SnapKV inevitably evicted tokens based on the first-turn conversation, it cannot successfully retrieve related information for future queries. In contrast, ShadowKV can maintain accuracy in the multi-turn conversation setting.
5.2 Efficiency Evaluation
To demonstrate the efficiency of ShadowKV, we deploy it into real-world large batch serving scenarios. By measuring the throughput during decoding across different models on A100, we show that ShadowKV can support up to larger batch sizes and boost throughput by up to 3.04.
Baselines.
The baseline selects the largest batch size that can fit entirely on the GPU with full attention. We also include results for the same batch size of ShadowKV and the infinite batch size, assuming infinite GPU memory capabilities444For the equivalent ShadowKV batch size, we evaluate a single Transformer block with FlashAttention and then project the number to the entire model. For the infinite batch size, we leverage A100’s theoretical memory bandwidth (2 TB/s) for attention computations.. We set the sparse budget to 1.56% for ShadowKV.
Results.
As shown in Table 4, ShadowKV demonstrates significant throughput improvements for various models on an A100, surpassing even those with infinite GPU memory. Notably, ShadowKV supports batch sizes up to larger and enhances throughput by up to compared to full attention, even surpassing infinite batch size assuming infinite GPU memory. While the gains for GLM-4-9B-1M and Yi-9B-200K are slightly lower, the improvements still reach up to and respectively, highlighting ShadowKV’s adaptability even with fewer KV heads.
Model | Context | Full Attention | ShadowKV | Gain | Full Attention (Inf) |
@@ 5.3 Ablation Results
We present extensive ablation studies of ShadowKV, focusing on three key points: (1) sparse KV cache budget variations, (2) chunk size selections, and (3) pre-RoPE key cache rank choices.
@@
As shown in Figure 9(a), increasing the chunk size allows for larger batch sizes. However, accuracy declines when the chunk size exceeds eight. Meanwhile, the chunk size choice has minimal impact on the chunk hit rate, which remains around 60%, as illustrated in Figure 9(b).
@@
We assess ShadowKV’s performance across various tasks using different ranks for pre-RoPE keys. As illustrated in Figure 9(c), accuracy increases with the rank up to approximately 160, after which it stabilizes near full-rank performance. Interestingly, the trends vary across tasks, and in some cases, low-rank approximations achieve better performance.
@@ 6 Conclusion
Appendix A Experiment Details
In this section, our goal is to provide the details of the system implementation (mentioned in Section 4.2), experiment settings, and additional experiments (mentioned in Section 5).
A.1 System Implementation.
We implement the framework based on PyTorch [36, 54] and dedicated kernels [50]. FlashAttention [8, 7, 18] is used for attention computation and some efficient fused kernels in Flashinfer [61] and vLLM [25] are used, including layer norm. To reduce memory movement and kernel launch overhead, we fuse some operations into CUDA kernels, including attention approximation, key cache low-rank reconstruction, value cache fetching, cache mechanism, etc. We leverage multi-streams to overlap the reconstruction of key cache and value cache fetching. We set the rank of pre-RoPE key cache to 160, chunk size to 8, and sparse KV cache budget to 1.56% for most cases.
A.2 Dataset Details
LLMs are widely used in various fields [55, 21, 28, 62, 39, 52, 44, 51, 17, 68, 6, 38, 37], and we select three long-context benchmarks, detailed below.
LongBench [4] is a challenging long-context benchmark that assesses the performance of LLMs in extended contexts. Featuring Chinese and English languages, LongBench encompasses 6 main categories and 21 diverse tasks, evaluating LLM capabilities across crucial long-text applications like single-/multi-document QA, summarization, code completion, etc.
Needle In A Haystack [24] is a long-context retrieval benchmark testing LLM’s performance with context window scales up to 1M tokens where information placed at various positions. We tested the retrieval capabilities of six long-context LLMs based on their context length.
A.3 Needle In A Haystack
In addition to the Needle In A Haystack results for Llama-3-8B-1M shown in Figure 6, we also present results for GLM-4-9B-1M, Llama-3.1-8B, Yi-9B-200K, Phi-3-Mini-128K, and Qwen2-7B-128K, shown in Figure 10. Compared to full attention, using ShadowKV has minimal impact on the ability to understand semantic information across different context windows and needle depths. There is even a slight performance improvement for Yi-9B-200K.
A.4 InfiniteBench
InfiniteBench [65] is a challenging long-context benchmark that consists of 10 tasks, including QA, coding, dialogue, summarization, and retrieval, with an average length of 214K.
Methods | En.Sum | En.QA | En.MC | En.Dia | Zh.QA | Code.Debug | Math.Find | Retr.PassKey | Retr.Num |
Llama-3-8B-1M | 23.05 | 18.14 | 65.06 | 10.50 | 12.47 | 24.36 | 37.14 | 100.00 | 100.00 |
ShadowKV | 21.50 | 17.73 | 64.63 | 10.50 | 12.45 | 23.86 | 37.43 | 100.00 | 100.00 |
GLM-4-9B-1M | 28.61 | 9.25 | 68.12 | 39.50 | 11.77 | 30.20 | 40.00 | 100.00 | 100.00 |
ShadowKV | 23.22 | 8.48 | 68.56 | 32.50 | 11.27 | 30.46 | 40.00 | 100.00 | 100.00 |
Llama-3.1-8B | 26.42 | 14.48 | 66.38 | 16.00 | 12.92 | 21.07 | 34.00 | 100.00 | 99.66 |
ShadowKV | 24.23 | 13.83 | 66.38 | 16.50 | 12.76 | 21.07 | 34.00 | 100.00 | 94.41 |
Yi-9B-200K | 8.88 | 10.61 | 61.57 | 5.50 | 13.88 | 21.57 | 23.71 | 100.00 | 99.66 |
ShadowKV | 8.92 | 10.06 | 59.39 | 6.00 | 13.89 | 20.56 | 24.29 | 100.00 | 99.83 |

### v2

@@ 4.2 Theoretical Equivalent Bandwidth
The benefit of ShadowKV in terms of increasing throughput can be analyzed through the concept of equivalent bandwidth. Consider each K or V vector as being bytes in size, with a sequence length of , a chunk size of , a selected chunk budget of , outliers, and hit rate . During KV selection, ShadowKV loads bytes using the GPU memory bandwidth . For value cache fetching, it loads bytes using the PCIe bandwidth [38]. Since value movement and key cache reconstruction can be overlapped, we do not need to count key cache reconstruction here. Following this, ShadowKV performs standard attention computation for the top-k chunks and predefined outliers, requiring bytes. The equivalent bandwidth of ShadowKV is defined as below and the GPU memory savings is detailed in Section 7.2.
For example, assuming C=8, S=128K, K=256, O=48, =31.5 GB/s, and =2 TB/s for A100, the equivalent bandwidth of ShadowKV is calculated as 7.2 TB/s, which is 3.6 higher than A100 memory bandwidth. This result indicates that ShadowKV theoretically achieves a high equivalent bandwidth to accelerate attention computation. System implementation is detailed in Section 8.1.
@@ 5 Empirical Evaluation
In Section 5.2, we demonstrate ShadowKV supports larger batch sizes and increase the inference throughput by up to without compromising model quality.
In Section 5.3, we present ablation studies that validate the effectiveness of each component of ShadowKV in optimizing GPU memory usage and performance.
@@ 5.1 Accuracy Evaluation
We demonstrate that ShadowKV can reduce GPU memory usage of the KV cache by while maintaining accuracy on a range of tasks with a minimal sparse KV budget.
@@
We choose four widely used long-context models for our evaluation: Llama-3-8B-1M [13], GLM-4-9B-1M [12], Llama-3.1-8B [31], and Yi-9B-200K [3]. We evaluate our approach on three challenging long-context benchmarks: RULER [17], LongBench [4], and Needle In A Haystack (NIAH) [20], covering QA, multi-hop, reasoning, summarization, code completion444We include results for Yi-9B-200K and other models (e.g., Llama-3-70B-1M) in Section 7. Needle In A Haystack is also tested on Phi-3-Mini-128K [1] and Qwen2-7B-128K [54].. We set the chunk size to 8, the rank to 160, and the number of outliers to 48 for ShadowKV.
@@
We include three dynamic sparse attention methods as baselines: Quest [45], Loki [39], and InfiniGen [23]. For all methods, we retain exact pre-filling and perform dynamic sparse attention during decoding, where the computation cost is set to 1/16 of full attention for selecting sparse KV pairs. We include two variants for each baseline: one where all the KV cache is offloaded, and another where only the value cache is offloaded. The former has similar latency to ShadowKV but a smaller sparse budget since ShadowKV only needs to fetch the value cache from the CPU. The latter aligns with the same sparse KV cache budget but significantly increases GPU memory usage. The latter one is marked as “(V)” in the table.
@@
As shown in Table 1, ShadowKV demonstrates excellent performance on 128K contexts. With a fixed sparse budget of 1.56%, other methods experience performance degradation. In contrast, ShadowKV is more robust and even outperforms original full attention on certain tasks, such as variable tracking. For complex tasks like multi-document QA or multi-key needle retrieval, other methods suffer from significant performance degradation while ShadowKV does not.
@@
On LongBench, we evaluate our method with a range of realistic scenarios, including single-/multi-document QA, document summarization, code completion, information retrieval, etc. We only test on samples longer than 4K and set the sparse KV cache budget to 256 for this benchmark since it has shorter inputs compared to RULER. As shown in Table 1, ShadowKV outperforms other methods consistently and maintains the performance.
NIAH.
On NIAH dataset, as shown in Figure 6, ShadowKV shows the ability to process information at different positions across various context windows, ranging from 16K to 1M tokens. More experiments on a range of models can be found in Section 8.3.
@@
Llama-3-8B-1M | 89.92 | 88.02 | 82.81 | 78.45 | 78.12 | 74.57 | 81.98 |
ShadowKV | 90.47 | 88.12 | 83.28 | 77.71 | 78.32 | 74.31 | 82.04 |
Multi-turn Capability.
To simulate multi-turn conversations, we challenged our method with a multi-turn needle retrieval benchmark (Multi-turn NIAH). We also test two eviction-based methods in Figure 7, including SnapKV [26] and StreamingLLM [52]. The performance of SnapKV drops significantly from the second round due to the required context information being different from the first round. Since SnapKV inevitably evicted tokens based on the first-turn conversation, it cannot successfully retrieve related information for future queries. In contrast, ShadowKV can maintain accuracy in the multi-turn conversation setting.
Model | Context | Full Attn | ShadowKV | Gain | Full Attn (Inf) |
@@
Context | 2 | 3 | 4 | 5 | 6 | 8 | 12 | 16 | 24 | 32 | 48 |
Full KV | |||||||||||
60K | 89.19 | 111.44 | 126.73 | 142.62 | 147.40 | 160.62 | OOM | OOM | OOM | OOM | OOM |
122K | 65.12 | 75.16 | 80.77 | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM |
244K | 40.37 | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM |
488K | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM | OOM |
ShadowKV | |||||||||||
60K | 89.69 | 126.61 | 159.41 | 184.92 | 205.41 | 244.20 | 306.09 | 346.34 | 399.65 | 428.63 | 455.14 |
122K | 65.61 | 94.28 | 115.01 | 132.23 | 143.77 | 166.72 | 196.73 | 217.24 | 239.51 | OOM | OOM |
244K | 48.39 | 65.95 | 78.92 | 87.83 | 94.07 | 104.73 | 119.01 | OOM | OOM | OOM | OOM |
488K | 29.82 | 41.01 | 47.13 | 50.85 | 53.46 | OOM | OOM | OOM | OOM | OOM | OOM |
5.2 Efficiency Evaluation
To demonstrate the efficiency of ShadowKV, we deploy it into real-world large batch serving scenarios. By measuring the throughput during decoding across different models on A100, we show that ShadowKV can support up to larger batch sizes and boost throughput by up to 3.04. The detailed latency breakdown can be found in Section 7.6.
Baselines.
The baseline selects the largest batch size that can fit entirely on the GPU with full attention. We also include results for the same batch size of ShadowKV and the infinite batch size, assuming infinite GPU memory capabilities555For the equivalent ShadowKV batch size, we evaluate a single Transformer block with FlashAttention and then project the number to the entire model. For the infinite batch size, we leverage A100’s theoretical memory bandwidth (2 TB/s) for attention computations.. We set the sparse budget to 1.56% for ShadowKV.
Results.
As shown in Table 3, ShadowKV demonstrates significant throughput improvements for various models on an A100, surpassing even those with infinite GPU memory. Notably, ShadowKV supports batch sizes up to larger and enhances throughput by up to compared to full attention, even surpassing infinite batch size assuming infinite GPU memory. While the gains for GLM-4-9B-1M and Yi-9B-200K are slightly lower, the improvements still reach up to and respectively, highlighting ShadowKV’s adaptability even with fewer KV heads. To further illustrate scalability, we provide detailed throughput measurements for Llama-3-8B-1M under varying batch sizes and context lengths in Table 4. As context length increases, full attention quickly runs out of memory, while ShadowKV continues to scale up to batch size 48 at 60K context and remains functional for extremely longer sequences. Additionally, we provide an efficiency comparison with Quest under 1M contexts in Section 7.7, demonstrating that ShadowKV significantly enhances throughput.
@@ 5.3 Ablation Results
In this section, we present extensive ablation studies of ShadowKV, focusing on three key points: (1) sparse KV cache budget variations, (2) chunk size selections, and (3) pre-RoPE key cache rank choices. Additional ablations, including precision sensitivity analysis and the effectiveness of outliers, are provided in Section 7.
@@
As shown in Figure 9, increasing the chunk size allows for larger batch sizes. However, accuracy declines when the chunk size exceeds eight. Meanwhile, the chunk size choice has minimal impact on the chunk hit rate, which remains around 60%.
@@
We assess ShadowKV’s performance across various tasks using different ranks for pre-RoPE keys. As illustrated in Figure 9, accuracy increases with the rank up to approximately 160, after which it stabilizes near full-rank performance. Interestingly, the trends vary across tasks, and in some cases, low-rank approximations achieve better performance.